The patient recovery time from a particular surgical procedure is normally distributed with a mean of 5.6 days and standard deviation of 1.9 days. Use your graphing calculator to answer the following questions. Write your answers in percent form. Round your answers to the nearest tenth of a percent.

a) What is the probability of spending less than 7 days in recovery? %

b) What is the probability of spending more than 5 days in recovery? %

c) What is the probability of spending between 5 days and 7 days in recovery? %
Answer #1

Solution :

Given that ,

mean = = 5.6

standard deviation = =1.9

a)

P(x < 7) = P[(x - ) / < (7 - 5.6) /1.9 ]

= P(z < 0.7368)

= 0.7694

probability = 7.69%

b)

P(x > 5) = 1 - P(x < 5)

= 1 - P((x - ) / < (5 - 5.6) / 1.9)

= 1 - P(z < -0.3158)

= 1 - 0.3761   

= 0.6239

Probability = 6.24%

c)

P(5 < x < 7) = P((5 - 5.6)/ 1.9) < (x - ) / < (7 - 5.6) / 1.9) )

= P(-0.3158 < z < 0.7368)

= P(z < 0.7368) - P(z < -0.3158)

= 0.7694 - 0.3761

= 0.3933

Probability = 3.93%